Operation Summary for March 26, 2010 - March 29, 2010 |
The Integrated Luminosity for the period from 3/22/10 to 3/29/10 was 61.3 pb-1. Friday March 26 The day shift began with Operations monitoring the Tevatron (TeV) store 7711 (233.57E30), the Recycler stash (8.1E10), and the Antiproton Source stack (41.7E10), with the Linac, Booster, and Main Injector operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations established beam to the Meson MTest experiment, Booster personnel reterminated a kicker thyratron cable (MKS011); this did not affect stacking or transfers to Recycler. During the evening shift, Operations reported that a Pbar sequencer was hanging up on a reverse proton command; operators rebooted a console. Operations disabled the MTest beam around 7 PM. The MI power supply interlock loop tripped off on a MI30 bypass loop trip; it recovered quickly. Saturday March 27 The midnight shift began with Operations monitoring store 7711 (62.58E30) and the stash (357.9E10), with Pbar stacking (25.2E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7711 at 2:55 AM and put the TeV into shot setup. Operations established store 7712 at 4:10 AM with an average initial luminosity (AIL) of 324E30. During the day shift, Operations reported that a Linac Klystron RF station (KRF3) tripped off as well as LRF1 and 3; everything reset. During the evening shift, Operations disabled the MTest beam at 6:01 PM. KRF3 tripped off twice and LRF3 tripped off once more. Operations terminated store 7712 at 10:23 PM and put the TeV into shot setup. Operations established store 7713 at 11:50 PM with an AIL of 345E30. Sunday March 28 The midnight shift began with Operations monitoring store 7713 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, the P1 transfer line beam permit tripped off on a vacuum burst. It took 20 minutes to recover vacuum and resume stacking. Operations established beam to MTest. During the day shift, Operations reported a TeV LCW pressure spike around sector A4; pressures returned to normal. During the evening shift, Operations terminated store 7713 at 5:35 PM. A States front end reboot had delayed the store termination and shot set up. Operations established store 7715 at 7:21 PM with an AIL of 358E30. Operations disabled beam to MTest. Monday March 28 The midnight shift began with Operations monitoring store 7715 and the stash, with Pbar stacking, with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to MTest. MIRF12 tripped off many times; operators contacted an RF person and then bypassed the station. MiniBooNE experimenters reported that they were having problems with the Dewatering system. Operators investigated and then called for advice. Controls personnel will investigate during the day shift. The Plans for Monday The plans for today are stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 03/29/2010 email Fermilab |
| Security, Privacy, Legal | |